### v2

# Softplus Attention with Re-weighting Boosts Length Extrapolation in Large Language Models

Abstract
Large language models have achieved remarkable success in recent years, primarily due to the implementation of self-attention mechanisms. However, traditional Softmax attention suffers from numerical instability and reduced performance as the length of inference tokens increases. This paper addresses these issues by decomposing the Softmax operation into a non-linear transformation and the $l_{1}$-norm. We identify the latter as essential for maintaining model performance. By replacing the non-linear transformation with the Softplus activation function and introducing a dynamic scale factor for different token lengths based on invariance entropy, we create a novel attention mechanism with performance better than conventional Softmax attention across various inference lengths. To further improve the length extrapolation ability of the proposed attention mechanism, we introduce a fine-tuning-free re-weighting mechanism that amplifies significant attention weights while diminishing weaker ones, enabling the model to concentrate more effectively on relevant tokens without requiring retraining. When combined with our proposed attention mechanism, this approach demonstrates significant promise in managing longer sequences, maintaining nearly constant validation loss even at 16$\times$ the training token length while ensuring numerical stability. Our code is available at: https://github.com/iminfine/freeatten.
Large Language Models; Softmax-free attention; Length Extrapolation
1 Introduction
Large Language Models (LLMs) have achieved significant success in recent years, largely due to the use of self-attention mechanism. This approach enables models to assess the importance of different words within a sentence, capturing complex relationships and dependencies in the data. Its effectiveness is mainly attributed to the Softmax operation. When the Softmax operation is omitted or replaced with alternative activation functions, model performance tends to decline (Wortsman et al., 2023; Ramapuram et al., 2024; Shen et al., 2023). Although widely used, the Softmax self-attention has two main limitations. Firstly, it suffers from numerical instability due to the natural exponential function, especially when scaling model sizes to trillions of parameters (sta, ). Secondly, as the token length increases during inference, the attention scores calculated by self-attention become smoother and lack distinct peaks. This hinders the model’s ability to establish connections between relevant tokens, thereby affecting the length extrapolation capabilities of transformers (Chiang & Cholak, 2022). This paper addresses these two problems simultaneously: developing a Softmax-free attention mechanism that offers improved numerical stability and better performance even at large token lengths.
Figure 1: Comparison of Softmax attention and the proposed re-weighted LSSA (LSSAR) across different inference sequence lengths.
Softmax calculates row attention scores in two steps: first, the scores are converted to a positive domain using the natural exponential function, and then each row is normalised to ensure the scores sum to one. This process inherently produces non-negative outputs, a trait long deemed vital for its success in attention mechanisms (Tsai et al., 2019; Katharopoulos et al., 2020; Han et al., 2023; Arora et al., 2024). However, this assumption has come under scrutiny, as studies substituting Softmax with other non-negative functions (Ramapuram et al., 2024; Li et al., 2022; Shen et al., 2023; Wortsman et al., 2023) report performance declines. Our experiments further explore this by assessing the necessity of non-negative attention scores, revealing that while non-negativity isn’t crucial, the normalisation process is key to sustaining performance.
Inspired by these observations, we decompose the Softmax operation into a non-linear transformation (the natural exponential function) and the $l_{1}$-norm, concluding that the $l_{1}$-norm is crucial to the success of traditional attention mechanisms. We then replace the non-linear transformation with various activation functions, resulting in models that match the performance of the original Softmax-based model. Furthermore, by substituting the non-linear transformation with the Softplus activation function and introducing a dynamic length scale factor that accounts for different token lengths based on invariance entropy, we propose a novel Length Scaled Softplus Attention (LSSA) mechanism. Our experiments demonstrate that LSSA outperforms the standard attention mechanism not only at the training sequence length but also for longer sequences. This highlights the potential of the proposed method to enhance the performance of LLMs.
To ensure that self-attention produces distinct peaks even for very long token length, we introduce a novel fine-tuning-free re-weighting approach that enhances the model’s capacity to concentrate on relevant tokens while filtering out noise. By applying a power transformation to the attention scores and $l_{1}$-norm afterwards, this re-weighting mechanism effectively amplifies stronger attention weights and suppresses weaker ones without retraining or architectural changes. Empirical results demonstrate that this re-weighting mechanism can be seamlessly integrated with various attention variants, leading to improved performance. When combined with our proposed LSSA, as shown in Figure 1 it shows particular promise in managing longer sequences, maintaining nearly constant validation loss even at 16 $\times$ the training token length while ensuring numerical stability.
In summary, this paper makes several important contributions:
  1. 1.
    We decompose the Softmax operation into a non-linear transformation and the $l_{1}$-norm and find the normalisation process is key to sustaining performance. By replacing the non-linear transformation with various activation functions, we develop multiple attention variants that are competitive with Softmax.
  2. 2.
    We introduce a novel Softmax-free attention mechanism, referred to as LSSA, which incorporates a dynamic length scale factor based on invariance entropy, demonstrating superior performance compared to standard attention mechanisms across a variety of sequence lengths.
  3. 3.
    We propose a novel fine-tuning-free re-weighting approach that enhances the model’s ability to focus on relevant tokens by amplifying significant attention weights while reducing weaker ones without requiring retraining or architectural modifications. When combined with our proposed LSSA, this re-weighted model (LSSAR) excels particularly in length extrapolation while ensuring numerical stability.
2 Method
2.1 Softmax Decomposition
Scaled dot-product attention transforms queries ($\mathbf{Q}$), keys ($\mathbf{K}$), and values ($\mathbf{V}$) through a two-step procedure. First, attention scores $\mathbf{A}$ are produced via:
$\mathbf{A}=\operatorname{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d}}+\mathbf{M}\right)$ (1)
Here, $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{L\times d}$, where $L$ represents the sequence length and $d$ denotes the dimensionality. The optional term $\mathbf{M}$ acts as a mask with shape $L\times L$ , which is particularly beneficial in causal self-attention contexts. The mask $\mathbf{M}$ is constructed with zeros below the diagonal and $\infty$- ∞ above it, ensuring that the attention mechanism only considers past and present tokens, thereby preserving the autoregressive nature of the model. The attention scores are subsequently utilised to compute the outputs as $\mathbf{O}=\mathbf{A}\mathbf{V}$.
Table 1: Validation loss values for training the GPT-2 model from scratch with the standard and modified attention scores.
Operation Softmax Inverse Re-centered Re-centered & $l_{1}$-norm
Loss 3.1911 3.1915 3.2008 3.1954
The Softmax function is a crucial component of the attention mechanism. It transforms a vector of raw scores $\mathbf{x}$ into a probability distribution, where the probabilities of each token sum to one. The outputs of Softmax are inherently non-negative, a feature considered essential for the success of Softmax attention. However, this assumption is somewhat questionable, as empirical studies that replaced the Softmax function with other non-negative output activation functions have shown a decline in model performance.
To investigate this further, we conducted experiments to test the necessity of non-negative attention scores. As illustrated in Table 1, one approach involved inverting the Softmax outputs by multiplying them by negative one, converting them to non-positive values. This inversion resulted in the validation loss remaining nearly unchanged, suggesting that non-negativity is not crucial for maintaining LLM performance. Additionally, we performed two more experiments. The first involved re-centring the attention scores. For each row of attention scores, we subtracted the mean value of that row, ensuring that each row had a mean of zero and, thus, contained both negative and positive values [^1] To prevent training instability due to the absence of positive values in the initial rows, the first three rows were left unmodified.. Inspired by the normalisation approach used in the Softmax operation, the second experiment further applied $l_{1}$-normalisation to each row, ensuring that the sum of the absolute values equaled one.
The results indicate that non-negative attention scores are not necessary for model performance. However, the $l_{1}$-norm is beneficial for maintaining performance. Motivated by these observations, we decomposed the Softmax operation into a non-linear transformation, the exponential function, followed by the $l_{1}$-norm:
$\operatorname{Softmax}(\mathbf{x})$ $=\frac{e^{\mathbf{x}}}{\sum_{j}e^{x_{j}}}=\frac{e^{\mathbf{x}}}{\|e^{\mathbf{x}}\|_{1}}=\frac{\phi(\mathbf{x})}{\|\phi(\mathbf{x})\|_{1}},$ (2)
[…]
Thus, the original attention function (Equation 1) can be expressed in a more general form:
$\mathbf{A}$ $=\phi\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d}}\right)\otimes\mathbf{M^{\prime}}$ (3)
$\mathbf{A}_{i}$ $\leftarrow\frac{\mathbf{A}_{i}}{\|\mathbf{A}_{i}\|_{1}}$
The subscripts $i$ and $j$ represent the row and column indices respectively. $\otimes$ denotes element-wise multiplication. Different from $\mathbf{M}$ defined in Equation 1, the mask $\mathbf{M^{\prime}}$ here is constructed with ones below the diagonal and zeros above it. It is evident that the original attention function (Equation 1) is a special case of Equation 3: one where the non-linear transformation function $\phi$ is the exponential function. This decomposition retains the essential characteristics of the attention mechanism while allowing for greater flexibility in the choice of activation functions. By replacing $\phi$ with various activation functions, we develop alternative attention mechanisms.
2.2 Length Scaled Softplus Attention
In this subsection, we present Length Scaled Softplus Attention (LSSA), a novel attention variant designed to enhance the scalability and performance of LLMs. Building upon the foundational work in attention mechanisms, LSSA introduces a novel scaling factor that accounts for both the sequence length and the model’s dimensionality, thereby addressing limitations associated with traditional attention methods in handling long sequences during inference.
LSSA utilises the Softplus activation function to introduce non-linearity while maintaining smooth gradients essential for stable training dynamics. Through empirical testing, we evaluated several widely-adopted activation functions by substituting them for $\phi$ into Equation 3. Our experiments in Section 3.2 revealed that Softplus outperforms the traditional exponential function ($e^{x}$), thereby justifying its adoption in the design of LSSA.
In contrast to scaled dot-product attention, cosine similarity attention employs the $l$2-norm for each row of $\mathbf{Q}$ and $\mathbf{K}$. This approach has been shown to produce more moderate attention weights, which can enhance performance across various tasks (Henry et al., 2020; Dehghani et al., 2023; Liu et al., 2022). However, in high-dimensional spaces, as the number of dimensions increases, two randomly selected vectors are likely to become orthogonal. This phenomenon causes the elements of the product of normalised $\|\mathbf{Q}\|_{2}$ and $\|\mathbf{K}\|_{2}^{T}$ to approach zero, thereby worsening the vanishing gradient problem. Consequently, a scale factor becomes essential for cosine similarity attention, and this factor should be associated with the dimensionality $d$.
Previous work demonstrated that replacing the traditional scaling factor $1/\sqrt{d}$ with $\log{L}/\sqrt{d}$ in scaled dot-product attention enhances the length extrapolation capabilities of transformers (Chiang & Cholak, 2022). Furthermore, (Su, 2021) highlighted that the inclusion of the $\log{L}$ factor aids in maintaining entropy invariance with different token length, thereby facilitating better extrapolation to unknown sequence lengths. Extending these findings, our approach adjusts the scaling factor to a dynamic length scale factor, reflecting the variation in token length across different rows of attention scores. Specifically, we set the scaling factor to $\log{d}\log{\mathbf{N}}$, where $\mathbf{N}$ is an $L\times L$ matrix, with each row containing identical elements corresponding to the number of tokens involved in attention for the $i$-th row, i.e., the number of tokens without masking. Clearly, $\mathbf{N}_{ij}=i$. This ensures that the attention mechanism remains robust across varying sequence lengths and model dimensions.
The formulation of LSSA is mathematically defined as:
$\mathbf{Q}_{i}$ $\leftarrow\frac{\mathbf{Q}_{i}}{\|\mathbf{Q}_{i}\|_{2}},\mathbf{K}_{i}\leftarrow\frac{\mathbf{K}_{i}}{\|\mathbf{K}_{i}\|_{2}}$ (4)
$\mathbf{A}$ $=\operatorname{Softplus}\left((\log{d}\log{\mathbf{N}})\otimes\mathbf{Q}\mathbf{K}^{T}\right)\otimes\mathbf{M^{\prime}}$
$\mathbf{A}_{i}$ $\leftarrow\frac{\mathbf{A}_{i}}{\|\mathbf{A}_{i}\|_{1}}$
It is important to note that due to hardware limitations, the proposed LSSA was evaluated with a sequence length of $L=1024$ and dimension $d=64$. For LLMs trained with longer sequence lengths and higher dimensionalities, the scaling factor may require further adjustment. However, the component $\log{\mathbf{N}}$ should remain unchanged, as it is crucial to maintain the length extrapolation capability of the model.
2.3 Attention Re-weighting Mechanism
While scaled dot-product attention has achieved remarkable success in LLMs, it inherently activates all input tokens to compute attention scores. This approach is somewhat counter-intuitive, as tokens positioned at the beginning of the input sequence are less likely to be relevant to tokens along the diagonal of row attention scores. Consequently, early tokens should ideally have minimal or no influence as $i$ increases when generating attention scores.
Experiments with ReLU-based attention, discussed in Section 3.2, demonstrate a minor performance decrease compared to the original attention mechanism. The observed performance decrement can be attributed to the distribution of positive values within the row attention scores. Specifically, positive values are predominantly concentrated around the diagonal, while other positions remain negative. The ReLU activation function tends to suppress these negative inputs, resulting in a limited number of activated neurons during training. This sparsity restricts the model’s capacity to capture essential relationships across the entire sequence, thereby adversely impacting overall performance. These results support the need for a re-weighting mechanism.
This phenomenon cannot be adequately addressed by traditional normalisation methods such as LayerNorm (Ba et al., 2016), as LayerNorm does not alter the underlying distribution of row attention scores. One potential solution is to transform all attention score values to the positive domain using functions like $e^{x}$ or Softplus before applying LayerNorm. As all values become positive through such transformations, employing the $l_{1}$-norm emerges as a more effective strategy. The $l_{1}$-norm is advantageous because it is simple to calculate and ensures that the mean value of the effective values of each row (the positive values that sum to one) remains constant (equals to $\frac{1}{i}$), thereby maintaining the integrity of the attention distribution. In this context, the Softmax operation in classic attention mechanism and Softplus with $l_{1}$-norm in LSSA can both be regarded as forms of normalisation layers and we can apply a re-weighting mechanism after them:
$\mathbf{A}$ $\leftarrow\operatorname{ReLU}^{p}\left(\mathbf{A}\otimes\mathbf{N}-\mathbf{O}\right)$ (5)
$\mathbf{A}_{i}$ $\leftarrow\frac{\mathbf{A}_{i}}{\|\mathbf{A}_{i}\|_{1}}$
where $\mathbf{O}$ denotes a matrix where every element is equal to one  [^2] To prevent training instability due to the absence of positive values in the initial rows, the elements within first three rows of $\mathbf{O}$ were set to zero.. The mean effective value of $\mathbf{A}\otimes\mathbf{N}$ is one, leading to an expectation of effective value of $\mathbf{A}\otimes\mathbf{N}-\mathbf{O}$ being zero. The notation $\operatorname{ReLU}^{p}$ indicates that the output of the ReLU function is raised to the power of $p$, with $p$ being a positive integer. ReLU effectively masks tokens with small attention scores, while the power operation enhances the discriminability of the remaining tokens. The re-weighting mechanism is applied post-training, requiring no additional gradient updates, which preserves the pre-trained model’s parameters intact and avoids the computational burden of fine-tuning.
As the token length increases, the attention scores become smoother and lacks distinct peaks, which hinders the model’s ability to establish connections between relevant tokens, thereby affecting the length extrapolation capabilities of transformers (Chiang & Cholak, 2022). This issue can be effectively addressed by the power operation, as demonstrated below:
Let $x_{1},x_{2},\ldots,x_{n}$ represent the positive elements in one row of the ReLU output from Equation 5, with the maximum value denoted as $x_{m}$. The formula for any input $x_{j}$ in Equation 5 is:
$\overline{x}_{j}=\frac{x_{j}^{p}}{\sum_{k=1}^{n}x_{k}^{p}}$ (6)
For any $x_{l}$ where $x_{l}<x_{m}$, the distance between their re-weighted values with respected to $p_{\to}\infty$ is given by:
$\lim_{p_{\to}\infty}{\Delta(\overline{x}_{m}-\overline{x}_{l})}$ $=\lim_{p_{\to}\infty}{\frac{x_{m}^{p}-x_{l}^{p}}{\sum_{k=1}^{n}x_{k}^{p}}}$ (7)
$=\lim_{p_{\to}\infty}{\frac{1-(\frac{x_{l}}{x_{m}})^{p}}{\sum_{k=1}^{n}(\frac{x_{k}}{x_{m}})^{p}}}$
$=1$
This limit demonstrates that as $p$ increases, the distance between the maximum value and any smaller value approaches 1. Specifically, when $p$ is large, the term $(\frac{x_{l}}{x_{m}})^{p}$ approaches zero for any $x_{l}<x_{m}$, causing all non-maximum values to be effectively suppressed. Meanwhile, the denominator becomes dominated by $x_{m}^{p}$, resulting in the maximum value approaching 1 while all other values approach 0. This property ensures that the attention mechanism maintains sharp, distinct peaks even with longer sequences, thereby preserving the model’s ability to focus on the most relevant tokens. We will refer to the combination of the re-weighting mechanism and LSSA as LSSAR in the rest of this paper.
3 Experiments
The experiments were executed using 8 NVIDIA A100 80GB GPUs. We utilised the GPT-2 small architecture, consisting of 124 million parameters (Radford et al., 2019). Importantly, we substituted the original absolute position embeddings with Rotary Position Embeddings (RoPE) (Su et al., 2024) and incorporated NTK Scaling (Liu et al., 2024b) to handle extended token lengths during inference for length extrapolation.
All models were trained using a sequence length of 1024 on the fineweb-10B dataset (Penedo et al., 2024), which consists of 10.2 billion training tokens distributed across 18,865 training steps, along with 0.1 billion validation tokens.
3.1 Softmax Decomposition
As illustrated in Equation 2, the Softmax operation can be decomposed into a non-linear transformation via the exponential function, followed by the $l_{1}$-norm. In this section, we examine the importance of each component for LLMs by training the GPT-2 model from scratch with and without each component.
Table 2: Validation loss for GPT-2 model with and without the Softmax components: $e^{x}$ and $l_{1}$-norm.
Scenario I II III IV
$e^{x}$
$l_{1}$-norm
Loss 3.4138 3.247 3.3297 3.1911
The results are presented in Table 2. Comparing the model employing Softmax in Scenario IV with Scenario I, which does not utilise Softmax, reveals that performance does not significantly deteriorate without Softmax. This can be attributed to the masking operation in attention, which elevates the rank of raw attention scores from $d$ to $L$. Given that $L\gg d$, causal masking substantially enhances the model’s non-linear capability. In tasks devoid of causal masking, such as transformers used in vision applications, performance degrades markedly (Wortsman et al., 2023). Scenario II shows that using a non-linearity enhances performance. However, it is possible that alternative activation functions may enhance performance beyond that produced by using the exponential function.
Moreover, the $l_{1}$-norm also contributes to performance enhancement. The attention score matrix constitutes a high-dimensional linear transformation matrix. By normalising each row of the transformation matrix to ensure that the sum of the absolute values equals one, the $l_{1}$-norm balances the influence of each dimension in the value vectors $\mathbf{V}$. This normalisation functions similarly to a weighted sum, permitting a more equitable contribution from each feature. Consequently, it stabilises the training process and prevents any single dimension from disproportionately impacting the output, ultimately leading to improved model performance.
3.2 Comparison with Different Activation Functions
We modify the standard attention mechanism by introducing several novel attention variants for comparative analysis. Specifically, these variants are created by substituting $\phi$ in Equation 3 with several widely-used activation functions: ReLU (Rumelhart et al., 1986), ReLU2 (So et al., 2021), ReLU6 (Howard et al., 2017), GeLU (Hendrycks & Gimpel, 2016), Sigmoid (Rumelhart et al., 1986), Softplus (Zheng et al., 2015), and Mish (Misra, 2019).
Table 3: Validation loss for various activation functions employed in the attention mechanism. Note that $\phi=e^{x}$ corresponds to the conventional Softmax attention.
ReLU ReLU2 ReLU6 GeLU Sigmoid Softplus Mish $e^{x}$
Loss 3.2006 3.2494 3.2039 3.2051 3.2000 3.1901 3.2001 3.1911
The results presented in Table 3 indicate that all activation functions, except for Softplus, result in higher validation loss values compared to the standard attention mechanism represented by $e^{x}$. This phenomenon can be attributed to the distribution of positive values in the row attention scores (the inputs to these activation functions), which are predominantly concentrated around the diagonal, while other positions remain negative. Consequently, these activation functions, apart from Softplus, tend to suppress negative inputs, leading to a limited number of activated neurons during training. This suppression results in higher loss values than those observed with the exponential function, $e^{x}$. Nevertheless, the performance differences are relatively minor, suggesting that the positive values play a more critical role than the negative values in influencing the model’s output.
Table 4: Comparison of state-of-the-art Softmax-free attention mechanisms across different sequence lengths. Bold font indicates the best result for each sequence length.
Attention Mechanism 1K 2K 4K 8K
Softmax 3.1911 4.1662 5.4513 6.2823
Sigmoid (Ramapuram et al., 2024) 3.1935 7.4554 11.8355 14.4995
Sigmoid (Ramapuram et al., 2024)($b=-\log{\mathbf{N}}$) 3.1930 6.5830 8.4679 9.5939
Sigmoid (Ramapuram et al., 2024)($b=-\log{\mathbf{N}},l_{1}$-norm) 3.1849 4.3470 5.5544 6.1846
Sigmoid (Proposed) 3.2000 4.3811 5.8465 6.5701
ReLU (Wortsman et al., 2023) 3.2143 6.2662 8.4982 10.3460
ReLU (Li et al., 2022) 3.2006 4.5192 5.6924 6.4561
ReLU (Shen et al., 2023) 3.2155 6.5573 8.9072 10.7266
LSSA 3.1905 4.1301 5.2960 5.9403
Re-weighted ($p=3$)
Softmax 3.1879 4.0277 5.2842 6.2339
Sigmoid (Ramapuram et al., 2024)($b=-\log{\mathbf{N}},l_{1}$-norm) 3.1841 3.7387 4.9286 5.7450
LSSAR 3.1782 4.2383 5.4056 6.3007
Re-weighted ($p=15$)
Softmax 5.3878 5.9491 6.5276 7.0183
Sigmoid (Ramapuram et al., 2024)($b=-\log{\mathbf{N}},l_{1}$-norm) 3.2171 3.3499 3.6108 3.8587
LSSAR 3.1905 3.1930 3.2291 3.3171
In contrast to $e^{x}$, the outputs of Softplus exhibit a slower growth rate, resembling a linear function when inputs are positive. This characteristic impedes its ability to emphasise large positive values, and consequently, the resulting attention focuses more on tokens that are further away from the diagonal, which could explain why its validation loss is lower than that of $e^{x}$. Further supporting this observation, the SquareReLU function clips negative values at zero and squares the positive inputs, thereby amplifying larger positive values even further, resulting in the highest loss value among the tested activation functions.
3.3 Comparison with State-of-the-Art Softmax-Free Attentions
We compared the proposed LSSA and LSSAR with leading Softmax-free attentions using the same GPT-2-124m model. These attentions include two Sigmoid attentions introduced by (Ramapuram et al., 2024) and the attention introduced in this paper (used in Table 3), and three ReLU attentions from (Wortsman et al., 2023), (Li et al., 2022), and (Shen et al., 2023).
For Sigmoid attention (Ramapuram et al., 2024), we evaluated two additional variants. The first variant alters the hyperparameter $b$ from $-\log{L}$ to $-\log{\mathbf{N}}$. The second variant maintains this modification while subsequently applying the $l_{1}$-norm to each row.
The experimental findings are detailed in Table 4. At an inference sequence length of 1K, all attention variants perform similarly, with only slight differences in validation loss. Specifically, Sigmoid(Ramapuram et al., 2024)($b=-\log{\mathbf{N}},l_{1}$-norm) and LSSA perform slightly better than the standard Softmax attention. As sequence length increases, performance differences become more evident. The standard Softmax attention experiences performance decline with longer sequences but remains relatively stable compared to most other methods.
Sigmoid attention (Ramapuram et al., 2024) shows significant performance decline at longer sequences, with validation loss rising sharply at 8K tokens. The modified Sigmoid(Ramapuram et al., 2024)($b=-\log{\mathbf{N}}$) shows some improvement, suggesting that the proposed hyperparameter $\log{\mathbf{N}}$ is more effective than $\log{L}$, although the model still faces challenges with longer sequences. Incorporating $l_{1}$-normalisation in Sigmoid(Ramapuram et al., 2024)($b=-\log{\mathbf{N}},l_{1}$-norm) aligns its performance more closely with standard Softmax attention, underscoring the importance of proper normalisation in attention mechanisms.
Figure 2: Comparison of LSSA (left figure) and Softmax attention (right figure) with varying values of $p$ for the re-weighting mechanism across different sequence lengths.
ReLU-based attention variants (Wortsman et al., 2023; Li et al., 2022) also experience notable performance decline with longer sequences, though not as drastically as the basic Sigmoid attention. This indicates that while ReLU-based methods may be effective for shorter sequences, they might not be ideal for managing longer-range dependencies.
To demonstrate the effectiveness of the proposed re-weighting mechanism, the attention scores generated by Softmax, Sigmoid (Ramapuram et al., 2024)($b=-\log{\mathbf{N}},l_{1}$-norm), and LSSA were re-weighted with $p=3$ and $p=15$. As shown in Table 4, the length extrapolation ability provided by the re-weighting can improve the performance of all the tested attention mechanisms. For Softmax performance was improved slightly using $p=3$, but was harmed when using $p=15$. For the other attention mechanisms $p=3$ was most effective for sequence lengths of 1K, while $p=15$ was best for the longer sequences. Notably, LSSAR exhibits exceptional stability across sequence lengths, maintaining nearly constant validation loss even at 8K tokens. This suggests that the re-weighting mechanism effectively addresses the challenges of processing longer sequences, possibly by better distributing attention across distant tokens.
3.4 Ablation Study for Re-weighting Mechanism
To gain deeper insights into the proposed re-weighting mechanism, we evaluated it using LSSA and Softmax attention across various values of $p$. As illustrated in Figure 2, LSSA and Softmax attention exhibit distinct responses to changes in the re-weighting parameter $p$. While Softmax attention performs comparably to LSSA at lower $p$ values, it suffers from severe gradient explosion issues as $p$ increases, particularly noticeable at $p=15$. This instability manifests as a sharp rise in validation loss across all sequence lengths, rendering the model practically unusable at higher $p$ values.
In contrast, increasing $p$ values generally improve the performance of LSSA across different sequence lengths, with optimal results observed around $p=15$. The validation loss remains relatively stable as the sequence length increases, underscoring LSSA’s strong scalability. This may be because transformers can be viewed as a smoothed cubic spline and Softplus activation is more natural, smooth, and maintains fidelity with spline representations (Lai et al., 2024). In addition, the derivative of the Softplus function is the Sigmoid function, which has an upper bound of one. Hence, LSSA is less susceptible to gradient explosion problems.
To further investigate LSSA’s stability with large $p$ values, we conducted additional tests with $p=50$ and $p=100$. LSSA maintains stability and avoids gradient explosion even at these extreme values, although we observe a gradual increase in validation loss. This increase occurs because larger $p$ values lead to a sharper attention distribution in each row, causing the model to focus primarily on the maximum value while potentially overlooking other relevant contextual information. This suggests that LSSA can effectively handle large $p$ values without numerical instability, outperforming Softmax attention in this aspect.
4 Discussion
While LSSAR was evaluated on a small model (GPT-2-124m) with a maximum inference token length of 16K, due to limited computing resources, we can infer from Figure 2 that the validation loss of LSSAR would remain relatively stable for token lengths greater than 16K. Furthermore the extreme $p$ value can be considered as increasing the depth of the GPT-2-124m model. In this context, it is reasonable to assume that LSSAR would also be stable for larger scale models with billions parameters.
The implementation of LSSA and LSSAR using PyTorch’s built-in functions incurs significant memory and computational costs. For practical applications, it is advisable to develop a highly optimised tensor program, such as Flashattention-3 (Shah et al., 2024), specifically tailored for these components. However, this requires a deep understanding of both hardware and CUDA programming, which can be a challenging task. As an alternative, it might be possible to utilise Mirage (Wu et al., 2024), a tool that automatically searches for suitable CUDA kernels and generates highly optimised tensor programs.
Additionally, it is recommended to use the T6 model (Zhang et al., 2025) with LSSAR in practical applications. T6 employs tensor decompositions to represent queries, keys, and values compactly, significantly reducing key-value cache memory while achieving better performance than standard transformers. By replacing Softmax attention with LSSAR in T6, the resulting model is expected to not only efficiently process significantly longer sequences under fixed resource constraints but also exhibit improved length extrapolation ability.
5 Related Work
Softmax-free Attention. Previous research has investigated Softmax-free attention by substituting the Softmax function with the ReLU activation (Li et al., 2022; Shen et al., 2023; Wortsman et al., 2023; Hron et al., 2020; Bai et al., 2024; Fu et al., 2024), SquaredReLU activation (Hua et al., 2022), and Sigmoid activation (Ramapuram et al., 2024), as well as examining purely linear attention (Qin et al., 2022; Tsai et al., 2019; Katharopoulos et al., 2020; Han et al., 2023; Arora et al., 2024; Lu et al., 2021). However, none of these approaches outperform the original Softmax attention. In contrast, the proposed LSSA demonstrates enhanced numerical stability and superior performance across various token lengths.
Attention Re-weighting. The non-linear re-weighting mechanism introduced by softmax attention ($l_{1}$-normalisation) has been shown to concentrate the distribution of attention weights, thereby stabilising the training process (AUEB et al., 2016; Gao & Pavel, 2017; Jang et al., 2016; Qin et al., 2022). Our empirical findings further demonstrate its essential role in maintaining the performance of LLMs. Moreover, we introduce a novel perspective: a non-linear positivity transformation followed by $l_{1}$-normalisation. Inspired by the classic normalisation-ReLU structure, the proposed re-weighting mechanism masks less relevant tokens and amplifies the relevant ones, which boosts the length extrapolation ability of underlying models. This provides the deep learning community with a deeper understanding of the attention mechanism within transformers.
Length Extrapolation. Positional embeddings (Su et al., 2024; Chen et al., 2023; Chi et al., 2022; Kiyono et al., 2021; Golovneva et al., 2024; He et al., 2024; Huang et al., 2020; Li et al., 2023; Likhomanenko et al., 2021; Liu et al., 2023; Wang et al., 2024; Zheng et al., 2024) play a vital role in transformer architectures by providing essential information about the positions of tokens within a sequence, which is considered a key factor in enhancing length extrapolation (Kazemnejad et al., 2023). Among these embeddings, RoPE is particularly noteworthy; it forms the foundation of many modern LLMs, including GPT-4 (Achiam et al., 2023), Llama3 (Dubey et al., 2024), Qwen2 (Yang et al., 2024), Deepseek-v3 (Liu et al., 2024a) and DeepSeek-R1 (DeepSeek-AI et al., 2025). While approaches like Position Interpolation (Chen et al., 2023; Li et al., 2023) modify positional encodings and require costly fine-tuning on extended sequences, the proposed LSSAR achieves robust length extrapolation without retraining, directly enhancing pre-trained models through its attention mechanism. This indicates that a well-designed attention mechanism like LSSAR is essential for RoPE-based LLMs to enhance their length extrapolation, suggesting that LSSAR may serve as a general strategy to bolster the length extrapolation abilities of most contemporary LLMs.
6 Conclusion
In this paper, we present a novel attention mechanism called LSSA. The primary innovation involves decomposing the Softmax operation, used in the traditional self-attention framework, into a non-linear transformation and the $l_{1}$-norm, replacing the non-linear transformation with the Softplus activation function and introducing a dynamic length scale factor related to token length. Experimental results indicate that LSSA surpasses conventional Softmax attention across various sequence lengths. To improve the length extrapolation capabilities of LSSA, we have devised a distinctive fine-tuning-free re-weighting mechanism that, when integrated with LSSA, the resulting model LSSAR significantly enhances its ability to extrapolate over longer sequences without requiring retraining, while maintaining stability and scalability, making it a practical, cost-effective solution for enhancing long-context capabilities in pre-trained LLMs without the computational burden of fine-tuning. This positions LSSAR as a compelling candidate for large-scale models. Consequently, we recommend the incorporation of LSSAR in future transformer architectures.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.